I have test corrections for homework. On the test, it says -
Write an equation of the line, in point-slope form, that is perpendicular to 3x+y=7 and passes through (-6,2).
I have done this on the test -
3x+y=7 (-3x on both sides)
y=-3x-7 (plug in coordinates)
y-2=-3(x+6)
My teacher said that the -3 part is wrong, yet I don't see why.

Please Reply Soon
Student in Need

Answer by lwsshak3(11628)

Write an equation of the line, in point-slope form, that is perpendicular to 3x+y=7 and passes through (-6,2).
**
Standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept
..
Given equation:
3x+y=7
y=-3x+7
slope=-3
..
Slope of line perpendicular to given equation=1/3 (negative reciprocal of -3)
y=(1/3)x+b=x/3+b
solving for b plugging in coordinates of given point (-6,2)
2=-6/3+b
2=-2+b
b=4
Equation of perpendicular line:
y=x/3+4